OpenStudy (jasmineflvs):

Given a = 2, b = -2, and k = 3, what is the equation of the graph if the parent function is y=square root of x ?

OpenStudy (anonymous):

does this come with a picture or some other equation with a, b, and k in it?

OpenStudy (jasmineflvs):

I will ask...

OpenStudy (jasmineflvs):

it's an equation on a graph...his computer won't let him copy and paste it

OpenStudy (anonymous):

hmmm... can he take a screenshot and paste that in?

OpenStudy (jasmineflvs):

he can't

OpenStudy (jasmineflvs):

can you get me an answer terenzreignz??

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

I just don't see any possible way without that graph, sorry.

terenzreignz (terenzreignz):

I don't know... what do a, b, and k even mean? D:

OpenStudy (calculusfunctions):

@Jasmineflvs a, b, and k would be parameters that are applied to the parent function. It depends on how they were defined in school. For example, with my students, I use a, k, d, and c where a represents the vertical dilation factor, k the horizontal dilation factor, d to describe the horizontal translation, and c to give the vertical translation.\[y=af[k(x -d)]+c\]For example if\[f(x)=2\sqrt{3(x +2)}\]where the parent function is\[f(x)=\sqrt{x}\]and a = 2, k = 3, d = -2, and c = 0. This would tell us that the graph of \[f(x)=\sqrt{x}\]is .vertically stretched by a factor of 2 .horizontally compressed by a factor of 1/3 .horizontally shifts 2 units left

OpenStudy (jasmineflvs):

so what is the answer...sorry I'm only helping a friend..he is almost in college...I am only in 10th grade

OpenStudy (jasmineflvs):

does anymore know it???

OpenStudy (calculusfunctions):

Well can your friend provide the parameters, that were used in class? The general form, like the one I provided above. It's not a matter of knowing it because I am a teacher, but you need to provide the the proper information, especially in such forums, in order for people to help.